If one were to use a standard normal distribution instead of Student's t distribution to calculate a confidence interval, what would be the "critical normal value" used as the coefficient for the s.e.m for a 90% CI?
Answer #1

Solution :

At 90% confidence level the z is ,

= 1 - 90% = 1 - 0.90 = 0.10

/ 2 = 0.10 / 2 = 0.05

Z/2 = Z 0.05 = 1.645

Critical normal value = 1.645